# Neural Gaffer: Relighting Any Object via Diffusion

### v1

@@ 1 Introduction
Lighting plays a key role in our visual world, serving as one of the fundamental elements that shape our interpretation and interaction with 3D space. The interplay of light and shadows can highlight textures, reveal contours, and enhance the perception of shape and form. In many cases, there is a desire to relight an image—that is, to modify it as if it were captured under different lighting conditions. Such a relighting capability enables photo enhancement (e.g., relighting portraits [16; 63]), facilitates consistent lighting across various scenes for filmmaking [55; 20], and supports the seamless integration of virtual objects into real-world environments [34; 17].
However, relighting single images is a challenging task because it involves the complex interplay between geometry, materials, and illumination. Hence, many classical model-based inverse rendering approaches aim to explicitly recover shape, material properties, and lighting from input images [1; 71; 43; 79; 83], so that they can then modify these components and rerender the image. These methods often require multi-view inputs and can suffer from: 1) model limitations that prevent faithful estimation of complex light, materials, and shape in real-world scenes, and 2) ambiguities in determining these factors without strong data-driven priors.
To circumvent these issues, image-based relighting techniques [16; 68; 72] avoid explicit scene reconstruction and focus on high-quality image synthesis. However, some image-based methods [16; 54; 72; 40; 2] require inputs under multiple lighting conditions or even sophisticated capture setups (e.g., a light stage [16]), while recent deep learning-based techniques [63; 47; 86] enable single-view relighting but often work only on specific categories like portraits.
@@ 2 Related Work
Diffusion models. Recently, diffusion-based image generation models [24; 45] have dominated visual content generation tasks. Stable Diffusion [56] not only generates impressive results in the text-to-image setting, but also enables a wide range of image applications. ControlNet [80] adds trainable neural network modules to the U-Net structure of diffusion models to generate images conditioned on edges, depth, human pose, etc. Diffusion models can also change other aspects of an input image [41; 8], such as image style, via image-to-image models [48]. ControlCom [76] and AnyDoor [12] are image composition models that can blend a foreground object into a target image. Huang et al. provide a comprehensive review of diffusion-based image editing [25].
@@
Recent work has also shown that, despite being trained on 2D images, other kinds of information can be unlocked from diffusion models. A notable example is Zero-1-to-3 [38], which uncovers 3D reasoning capabilities in Stable Diffusion by fine-tuning it on a view synthesis task using synthetic data, and conditioning the model on viewpoint. Similar to the spirit of Zero-1-to-3 and its follow-up work (such as [59]), our work unlocks the latent capability of diffusion models for relighting.
@@
Single-image relighting. Relighting imagery is a difficult task that requires an (explicit or implicit) understanding of geometry, materials, lighting, and their interaction through light transport. Single-image relighting is especially challenging due to its ill-posedness; there are infinite explanations of a single image in terms of underlying factors. Classic single-image methods, like SIRFS [1], focus on explicit estimation of scene properties, via optimization with hand-crafted priors. More recent learning-based methods can learn such priors from data, but either focus on specific image types, such as outdoor scenes [69; 65; 74; 36; 19], portraits [49; 67; 60; 63; 44; 47; 21; 52], or human bodies [28; 33; 26], or require special capture conditions like camera flash [35; 57]. Single-image relighting methods for general objects, with realistic appearance, materials, and lighting, remain elusive. In addition, most recent single-image relighting methods [47; 35; 57] still explicitly decompose an image into components, then reconstruct it with modified lighting. This strategy allows the components to be explicitly controlled, but also limits their realism based on the representational ability of the chosen components, such as the Disney BSDF [10] or Phong reflection model [9]. In contrast, similar to [63; 75], we achieve object relighting without an explicit scene decomposition. By training an image-based and lighting-conditioned diffusion model, we can relight any given object from a single photo.
@@
3D relighting and inverse rendering. Inverse rendering methods involve estimating 3D models that can be relit under different lighting conditions, typically from multiple images. Recent advances in 3D, such as NeRF [42] and 3D Gaussian Splatting [29], have shown remarkable novel view synthesis results. However, these representations are not themselves relightable as the physical light transport is baked into the radiance field. Efforts have been made toward creating relightable 3D representations, such as PhySG [79], NeRFactor [83], TensoIR [27] and others [5; 7; 6; 4; 3; 78; 84; 31; 62; 73; 77]. These methods often rely on pure optimization to minimize a rendering loss, sometimes incorporating hand-crafted material and/or lighting regularization terms (e.g., smoothness or sparsity) to prevent overfitting. However, these approaches often struggle with specular highlights and glossy materials. We show that our proposed Neural Gaffer method can be applied in an inverse rendering framework, serving as a powerful prior and regularizer. This is because our model has been trained on a large dataset encompassing diverse lighting and material conditions, effectively learning these priors from high-quality data. Our model improves 3D relighting quality, while baselines tend to produce artifacts due to the baking of lighting information into the color field.
@@ 3.1 RelitObjaverse Dataset
Capturing ground-truth lighting for real object images is both time-consuming and challenging. Existing real-world relighting datasets [37; 32; 66] are constrained to a very limited number of objects. To address this limitation, we opt to render a large-scale synthetic object relighting dataset.
@@
We render images with the Cycles renderer from Blender [15]. During rendering, we randomly sample $12$ camera poses for each object. For each viewpoint, we render $16$ images under different environment maps. In contrast to existing datasets that only render with environment maps [14; 58], we additionally render an image under a randomly set area light source to increase lighting diversity. In total, for each object, there are ${12\times(16+1)=204}$ images rendered under different lighting conditions and from different viewpoints. Our final synthetic relighting dataset has about 18.4M ${512\times 512}$ rendered images with ground-truth lighting. We will release our full relighting dataset, named the RelitObjaverse dataset, upon acceptance to facilitate reproducibility.
@@ 3.2 Lighting-conditioned 2D Relighting Diffusion Model
HDR-LDR conditioning. Pixel values in the HDR environment map have a theoretical range of $[0,+\infty)$, and large pixel values can lead to unstable diffusion training. To rescale it to $[0,1.0]$, a naive idea is to divide it by its maximum value, but this leads to a loss of total energy information. Also, if the map has a very bright region (e.g., the sun), then most of the other values will approach 0 and the environment map will be too dark after division, causing a loss of lighting detail. To solve this, we split the HDR environment map $E$ into two maps: $E_{L}$ is an LDR representation of $E$, computed via standard tone-mapping [53; 22], and $E_{H}$ is a nonlinearly normalized representation of $E$, computed by a logarithmic mapping and divided by the new maximum value. $E_{L}$ encodes lighting detail in low-intensity regions, which is useful for generating relighting details(such as reflection details). $E_{H}$ maintains the lighting information across the initial spectrum. Inputting $E_{L}$ and $E_{H}$ together to the diffusion model means that the network can reason effectively about the full original energy spectrum. Additionally, our design helps to achieve a balanced exposure in the generated output, preventing areas from appearing too dark or too washed out after normalization due to extreme brightness.
@@ 3.3 Relighting a 3D radiance field with diffusion prior guidance
Stage 2: Detail refinement with diffusion guidance loss. The results from the first stage can be blurry and can exhibit noise because the multi-view relighting predictions results may be inconsistent. We further refine the relit NeRF’s appearance in Stage 2. We found that directly using the SDS loss [51] for refinement leads to noisy and over-saturated results (see Fig.  8). Inspired by [64; 70; 41], we treat the blurry rendered images from Stage 1 as an intermediate denoising result. We first render an image $\hat{x}$ under the sampled camera pose $\pi_{\text{sampled}}$. We encode and perturb it to a noisy latent $\hat{z_{t}}(\hat{x},t^{\star})$ with an intermediate noise level $t^{\star}$, and initiate the multi-step DDIM [61] denoising process for $\hat{z_{t}}(\hat{x},t^{\star})$ from $t^{\star}$, yielding a latent sample $\hat{z}_{0}$. This latent is decoded to an image $\hat{x}_{\text{relit}}(t^{\star})=D(\hat{z_{0}})$, which we use to supervise the rendering with the following diffusion guidance loss:
@@ 4.2 Baselines and metrics
We use common image metrics including PNSR, SSIM, and LPIPS [82] to evaluate relighting results quantitatively. In addition, since there exists an ambiguity between lighting intensity and object color, we also compute channel-aligned results as in [83; 85], where each channel of the relighting prediction is rescaled to align the average intensity with the ground-truth. We compute metrics for both the raw and the channel-aligned results for quantitative evaluation.
@@ 4.3 Results analysis
Relighting 3D objects. Fig. 6 shows the results of our two-stage method for relighting 3D objects. Given an input radiance field(NeRF) representing a 3D object, we first tune NeRF for coarse relighting in Stage 1, followed by a finer relighting result in Stage 2 that improves fidelity. We also compared to conventional inverse rendering methods such as NVDIFFREC-MC [23] and TensoIR [27], reporting quantitative results in Tab. 2 and showing qualitative comparisons in Fig. 6.
@@ 5 Limitations
While our model can achieve overall consistent relighting results under rotating or changing lighting conditions, our results can still exhibit minor inconsistencies since the model is generative. Due to the limited resolution of the backbone diffusion model, we can only handle relatively low image resolution ($256\times 256$). In addition, our model is fine-tuned on object data, therefore, while it is category-agnostic, it might not achieve the highest quality results on specific domains—for instance, for the case of portrait relighting, methods that are specifically trained for portrait photos may produce higher quality results.
@@ 6 Conclusion
We proposed an end-to-end 2D relighting diffusion model, Neural Gaffer, that addresses the long-standing challenge of single-image relighting through a data-driven framework. Our model synthesizes accurate and high-quality relighted images from a single input image under any environmental lighting condition, demonstrating superior generalization and accuracy on both synthetic and real-world datasets. Neural Gaffer not only enhances various 2D tasks, such as text-based relighting and object insertion, but also serves as a strong relighting prior for 3D tasks, enabling a simplified two-stage 3D relighting pipeline. By overcoming the limitations and ambiguities of traditional model-based inverse rendering approaches, our diffusion-based solution extends the applicability of relighting techniques to a broad range of practical applications.
@@
This work was done while Haian Jin was a full-time student at Cornell. The selection of data and the generation of all figures and results was led by Cornell University. This work was funded in part by the National Science Foundation (IIS-2211259).

### v2

@@ 1 Introduction
Lighting plays a key role in our visual world, serving as one of the fundamental elements that shape our interpretation and interaction with 3D space. The interplay of light and shadows can highlight textures, reveal contours, and enhance the perception of shape and form. In many cases, there is a desire to relight an image—that is, to modify it as if it were captured under different lighting conditions. Such a relighting capability enables photo enhancement (e.g., relighting portraits [18; 67]), facilitates consistent lighting across various scenes for filmmaking [59; 22], and supports the seamless integration of virtual objects into real-world environments [38; 19].
However, relighting single images is a challenging task because it involves the complex interplay between geometry, materials, and illumination. Hence, many classical model-based inverse rendering approaches aim to explicitly recover shape, material properties, and lighting from input images [2; 75; 47; 83; 87], so that they can then modify these components and rerender the image. These methods often require multi-view inputs and can suffer from: 1) model limitations that prevent faithful estimation of complex light, materials, and shape in real-world scenes, and 2) ambiguities in determining these factors without strong data-driven priors.
To circumvent these issues, image-based relighting techniques [18; 72; 76] avoid explicit scene reconstruction and focus on high-quality image synthesis. However, some image-based methods [18; 58; 76; 44; 3] require inputs under multiple lighting conditions or even sophisticated capture setups (e.g., a light stage [18]), while recent deep learning-based techniques [67; 51; 89] enable single-view relighting but often work only on specific categories like portraits.
@@ 2 Related Work
Diffusion models. Recently, diffusion-based image generation models [26; 49] have dominated visual content generation tasks. Stable Diffusion [60] not only generates impressive results in the text-to-image setting, but also enables a wide range of image applications. ControlNet [84] adds trainable neural network modules to the U-Net structure of diffusion models to generate images conditioned on edges, depth, human pose, etc. Diffusion models can also change other aspects of an input image [45; 9], such as image style, via image-to-image models [52]. ControlCom [80] and AnyDoor [14] are image composition models that can blend a foreground object into a target image. Huang et al. provide a comprehensive review of diffusion-based image editing [28].
@@
Recent work has also shown that, despite being trained on 2D images, other kinds of information can be unlocked from diffusion models. A notable example is Zero-1-to-3 [42], which uncovers 3D reasoning capabilities in Stable Diffusion by fine-tuning it on a view synthesis task using synthetic data, and conditioning the model on viewpoint. Similar to the spirit of Zero-1-to-3 and its follow-up work (such as [63; 27]), our work unlocks the latent capability of diffusion models for relighting.
@@
Single-image relighting. Relighting imagery is a difficult task that requires an (explicit or implicit) understanding of geometry, materials, lighting, and their interaction through light transport. Single-image relighting is especially challenging due to its ill-posedness; there are infinite explanations of a single image in terms of underlying factors. Classic single-image methods, like SIRFS [2], focus on explicit estimation of scene properties, via optimization with hand-crafted priors. More recent learning-based methods can learn such priors from data, but either focus on specific image types, such as outdoor scenes [73; 69; 78; 40; 21], portraits [53; 71; 64; 67; 48; 51; 23; 56], or human bodies [31; 37; 29], or require special capture conditions like camera flash [39; 61]. Single-image relighting methods for general objects, with realistic appearance, materials, and lighting, remain elusive. In addition, most recent single-image relighting methods [51; 39; 61; 12] still explicitly decompose an image into components, then reconstruct it with modified lighting. This strategy allows the components to be explicitly controlled, but also limits their realism based on the representational ability of the chosen components, such as the Disney BSDF [11] or Phong reflection model [10]. In contrast, similar to [67; 79], we achieve object relighting without an explicit scene decomposition. By training an image-based and lighting-conditioned diffusion model, we can relight any given object from a single photo.
@@
3D relighting and inverse rendering. Inverse rendering methods involve estimating 3D models that can be relit under different lighting conditions, typically from multiple images. Recent advances in 3D, such as NeRF [46] and 3D Gaussian Splatting [32], have shown remarkable novel view synthesis results. However, these representations are not themselves relightable as the physical light transport is baked into the radiance field. Efforts have been made toward creating relightable 3D representations, such as PhySG [83], NeRFactor [87], TensoIR [30] and others [6; 8; 7; 5; 4; 82; 88; 35; 66; 77; 81]. These methods often rely on pure optimization to minimize a rendering loss, sometimes incorporating hand-crafted material and/or lighting regularization terms (e.g., smoothness or sparsity) to prevent overfitting. However, these approaches often struggle with specular highlights and glossy materials. We show that our proposed Neural Gaffer method can be applied in an inverse rendering framework, serving as a powerful prior and regularizer. This is because our model has been trained on a large dataset encompassing diverse lighting and material conditions, effectively learning these priors from high-quality data. Our model improves 3D relighting quality, while baselines tend to produce artifacts due to the baking of lighting information into the color field.
@@ 3.1 RelitObjaverse Dataset
Capturing ground-truth lighting for real object images is both time-consuming and challenging. Existing real-world relighting datasets [41; 36; 70] are constrained to a very limited number of objects. To address this limitation, we opt to render a large-scale synthetic object relighting dataset.
@@
We render images with the Cycles renderer from Blender [17]. During rendering, we randomly sample $12$ camera poses for each object. For each viewpoint, we render $16$ images under different environment maps. In contrast to existing datasets that only render with environment maps [16; 62], we additionally render an image under a randomly set area light source to increase lighting diversity. In total, for each object, there are ${12\times(16+1)=204}$ images rendered under different lighting conditions and from different viewpoints. Our final synthetic relighting dataset has about 18.4M ${512\times 512}$ rendered images with ground-truth lighting. We will release our full relighting dataset, named the RelitObjaverse dataset, upon acceptance to facilitate reproducibility.
@@ 3.2 Lighting-conditioned 2D Relighting Diffusion Model
HDR-LDR conditioning. Pixel values in the HDR environment map have a theoretical range of $[0,+\infty)$, and large pixel values can lead to unstable diffusion training. To rescale it to $[0,1.0]$, a naive idea is to divide it by its maximum value, but this leads to a loss of total energy information. Also, if the map has a very bright region (e.g., the sun), then most of the other values will approach 0 and the environment map will be too dark after division, causing a loss of lighting detail. To solve this, we split the HDR environment map $E$ into two maps: $E_{L}$ is an LDR representation of $E$, computed via standard tone-mapping [57; 24], and $E_{H}$ is a nonlinearly normalized representation of $E$, computed by a logarithmic mapping and divided by the new maximum value. $E_{L}$ encodes lighting detail in low-intensity regions, which is useful for generating relighting details(such as reflection details). $E_{H}$ maintains the lighting information across the initial spectrum. Inputting $E_{L}$ and $E_{H}$ together to the diffusion model means that the network can reason effectively about the full original energy spectrum. Additionally, our design helps to achieve a balanced exposure in the generated output, preventing areas from appearing too dark or too washed out after normalization due to extreme brightness.
@@ 3.3 Relighting a 3D radiance field with diffusion prior guidance
Stage 2: Detail refinement with diffusion guidance loss. The results from the first stage can be blurry and can exhibit noise because the multi-view relighting predictions results may be inconsistent. We further refine the relit NeRF’s appearance in Stage 2. We found that directly using the SDS loss [55] for refinement leads to noisy and over-saturated results (see Fig.  8). Inspired by [68; 74; 45], we treat the blurry rendered images from Stage 1 as an intermediate denoising result. We first render an image $\hat{x}$ under the sampled camera pose $\pi_{\text{sampled}}$. We encode and perturb it to a noisy latent $\hat{z_{t}}(\hat{x},t^{\star})$ with an intermediate noise level $t^{\star}$, and initiate the multi-step DDIM [65] denoising process for $\hat{z_{t}}(\hat{x},t^{\star})$ from $t^{\star}$, yielding a latent sample $\hat{z}_{0}$. This latent is decoded to an image $\hat{x}_{\text{relit}}(t^{\star})=D(\hat{z_{0}})$, which we use to supervise the rendering with the following diffusion guidance loss:
@@ 4.2 Baselines and metrics
We use common image metrics including PNSR, SSIM, and LPIPS [86] to evaluate relighting results quantitatively. In addition, since there exists an ambiguity between lighting intensity and object color, we also compute channel-aligned results as in [87; 88], where each channel of the relighting prediction is rescaled to align the average intensity with the ground-truth. We compute metrics for both the raw and the channel-aligned results for quantitative evaluation.
@@ 4.3 Results analysis
Relighting 3D objects. Fig. 6 shows the results of our two-stage method for relighting 3D objects. Given an input radiance field(NeRF) representing a 3D object, we first tune NeRF for coarse relighting in Stage 1, followed by a finer relighting result in Stage 2 that improves fidelity. We also compared to conventional inverse rendering methods such as NVDIFFREC-MC [25] and TensoIR [30], reporting quantitative results in Tab. 2 and showing qualitative comparisons in Fig. 6 and Fig. 13.
@@ 5 Limitations
Figure 9: Limitation Analysis. The left side shows the encoding-decoding results using the VAE [33] in our base diffusion model [1; 42]. These results indicate that the VAE struggles to preserve the identity of objects with fine details at low resolution ($256\times 256$), which in turn affects the performance of our model on such objects. On the right side, we present several failed cases of our method in Objaverse-like instances.
While our model can achieve overall consistent relighting results under rotating or changing lighting conditions, our results can still exhibit minor inconsistencies since the model is generative.
Due to the significant computational demands of data preprocessing (specifically, rotating the HDR environment map) and model training, and considering our limited university resources, we trained the model at a low image resolution ($256\times 256$). This choice has become a major limitation in our work. As illustrated in Fig. 9, we use the VAE [33] in our base diffusion model [1; 42] to encode input images into latent maps, and then directly decode them. We observed that the VAE struggles to preserve the identity of objects with fine details, even when the decoded latent maps are directly derived from input images at this resolution. Consequently, this leads to numerous relighting failures. Fine-tuning our model at a higher resolution would significantly mitigate these issues.
We observe two key failure cases when relighting Objaverse-like instances, as shown on the right side of Fig. 9. First, our model struggles with color ambiguity in diffuse objects, leading to inaccuracies in relit colors (first row). This issue arises from inherent ambiguities in the input image, where color could be attributed to either material or lighting. Including the background of the input image may help the model better infer the lighting conditions of input images and address this challenge. We give more discussions in the Appendix. C. Second, while the model generates low-frequency highlights well, it often misses high-frequency reflections in relit images (second row).
In addition, we discuss our limitations regarding portrait relighting in the Appendix. B.
@@ 6 Conclusion
We proposed an end-to-end 2D relighting diffusion model, Neural Gaffer, that addresses the long-standing challenge of single-image relighting through a data-driven framework. Our model synthesizes accurate and high-quality relighted images from a single input image under any environmental lighting condition, demonstrating superior generalization and accuracy on both synthetic and real-world datasets. Neural Gaffer not only enhances various 2D tasks, such as text-based relighting and object insertion but also serves as a strong relighting prior for 3D tasks, enabling a simplified two-stage 3D relighting pipeline. By overcoming the limitations and ambiguities of traditional model-based inverse rendering approaches, our diffusion-based solution extends the applicability of relighting techniques to a broad range of practical applications.
@@
This work was done while Haian Jin was a full-time student at Cornell. The selection of data and the generation of all figures and results was led by Cornell University. This work was funded in part by the National Science Foundation (IIS-2211259). Jin Sun is partly supported by a gift from Google.